In the following ordered data, the mean is 6 and the median is 5.

2,b, 3, a, 6, 9, 10, 12 Find each of the following (a) the value of a;

(b) the value of b;


1
Expert's answer
2021-11-26T13:51:40-0500
"mean=\dfrac{2+b+3+a+6+9+10+12}{8}=6"

"a+b=6"

"median=\dfrac{a+6}{2}=5"

(a)


"a=5(2)-6"

"a=4"

(b)


"4+b=6"

"b=2"